OpenStudy (anonymous):

find the domain of the function f(x)= 1/sqrt (4x-8) please show calculations

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

callum can you show the steps please

OpenStudy (anonymous):

(2, infinity)

OpenStudy (anonymous):

\[f(x) = \frac{1}{\sqrt{4x-8}}\] now you want to look at conditions on 4x-8. What can x reasonably be? Well you cannot take a square root of a negative number, and the denominator cannot be zero, so we need: \[4x-8 > 0\] \[x>\frac{1}{2}\] So I was wrong above, the domain is actually: \[\left(\frac{1}{2}, \infty \right)\]

OpenStudy (anonymous):

x > 2

OpenStudy (anonymous):

think you divided the wrong side there

OpenStudy (anonymous):

oh yes sorry it's 2 lol, quite right. \[(2,\infty)\]

OpenStudy (anonymous):

rickjbr can you demonstrate how you got your answer please

OpenStudy (anonymous):

you'll have imaginary numbers if the square root is allowed to go negative

OpenStudy (anonymous):

thus not part of the real number system

OpenStudy (anonymous):

callum was right in his calculation 4x - 8 > 0 4(x-2) > 0 x-2 > 0 x > 2

OpenStudy (anonymous):

yeah my reasoning was right, it's just I divided the inequality wrong and got x>1/2. rickjbr corrected this to x>2

OpenStudy (anonymous):

thanks for the help to both of your

OpenStudy (anonymous):

np

OpenStudy (anonymous):

how about this. 2x^2-3x+5...find and simplify the difference quotient

OpenStudy (anonymous):

The question reads Find and simplify the difference quotient f(x+h)-f(x)/h for the given function f(x)=2x^(2)-3x+5

OpenStudy (anonymous):

ah, right

OpenStudy (anonymous):

(2(x+h)^2-3(x+h)+5-2x^2+3x-5)/h

OpenStudy (anonymous):

(2(x^2+2xh+h^2)-3(x+h)+5-2x^2+3x-5)/h

OpenStudy (anonymous):

(2x^2+4xh+2h^2-3x-3h+5-2x^2+3x-5)/h

OpenStudy (anonymous):

(2h^2+4xh+3x)/h

OpenStudy (anonymous):

is that the final answer

OpenStudy (anonymous):

checking, a bit rusty

OpenStudy (anonymous):

my options are a)4x+2h, b)4x+2h-3, c)4h+2x-3, d)4h+2x, e)4x+4h

OpenStudy (anonymous):

i'll redo it, might've missed something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ah, silly

OpenStudy (anonymous):

that last step was it. instead of (2h^2+4xh+3x)/h it should be (2h^2+4xh-3h)/h

OpenStudy (anonymous):

simply factor out an h

OpenStudy (anonymous):

[h(2h+4x-3)]/h =2h+4x-3

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

thanks greatly. i dont know if you have the time...but im reviewing for a test and I have a whole lot of other questions. if you can't help me, i understand, i will just post on the main forum

OpenStudy (anonymous):

probably better off posting on main forum, if i can help, i'll try

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

Find \[\frac{f(x+h)-f(x)}{h}\] for \[f(x)=2x^2-3x+5\] \[\frac{f(x+h)-f(x)}{h} = \frac{(2(x+h)^2-3(x+h)+5) - (2x^2-3x+5)}{h}\] \[= \frac{(2x^2+4xh+2h^2-3x-3h)-(2x^2-3x)}{h}\] \[=\frac{4xh+2h^2-3h}{h} = 4x+2h-3\] We can verify that this is right by noting that \[f'(x) = \lim_{h\rightarrow 0}\frac{f(x+h) - f(x)}{h}\] The left hand side of this is equal to the derivative of f, which we can just do right here: \[f'(x) = 4x - 3\] and we should get that this equals the limit as h goes to zero of \[\frac{f(x+h) - f(x)}{h}.\] We have already calculated that \[\frac{f(x+h) - f(x)}{h} = 4x+2h-3\] so taking the limit of this as h goes to zero we get: \[\lim_{h\rightarrow 0}(4x+2h - 3) = 4x+0-3 = 4x-3.\] Hence we ave verified that the simplification has been done correctly.